A rectangular garden is 42 feet longer than it is wide its perimeter is 312 feet write a system of linear equations whose solution gives the width and length of the garden use a method of substitution to solve the system
Answer by Cromlix(4381)

Hi there,
Make the garden's width = 'x'
Length of garden = x + 42
Perimeter = 2(Width) + 2(Length)
312 = 2(x) + 2(x + 42)
312 = 2x + 2x + 84
Collect like terms
4x = 312 - 84
4x = 228
x = 57
Width = 57 feet
Length = 99 feet
Hope this helps :-)